OpenStudy (anonymous):

y = sqrt. of (36 - x^2)

OpenStudy (anonymous):

What are you meant to do with it? You can't really simplify it, unless you want to use the fact that 36-x^2 is a difference of two squares and expand the square root as\[\sqrt{36-x^2}=\sqrt{(6-x)(6+x)}=\sqrt{6-x}\sqrt{6+x}\]

OpenStudy (anonymous):

If you're not using complex numbers than it's just solvable for: \[-6<x<6\] For all other cases the term under the root is less than zero, which is not defined.